OpenStudy (anonymous):

p=x-2 q=-x^2+2 find: (p*q)(2)= (q*p)(2)=

OpenStudy (anonymous):

just something to go off of

OpenStudy (anonymous):

Assuming that the sign * is for (of) not multiplication, you can follow this: \[p(x)=x-2\]\[p(2)=2-2\]\[p(2)=0\] \[q(x)=x ^{2}+2\]\[q(2)=2 ^{2}+2\]\[q(2)=4+2\]\[q(2)=6\] So, (p*q)(2)= p(q(2))= 6-2=4 (q*p)(2)= q(p(2))= 0^2+2=2. If the sign * is multiplication then. (p*q)(2)= (0)(6)=0 (q*p)(2)=(6)(0)=0 Hope this helps.

OpenStudy (anonymous):

the * is for multiplication!

OpenStudy (anonymous):

awesome man thanks that helps me to so i might be able to help someone in the future